# TIT/HEP-708

### v1

@@ 4 TBA for the cubic superpotential
In this section, we illustrate our general procedure of wall-crossing with the case of cubic superpotential, where the Schrödinger equation (1) becomes
@@
where $k^{2}=1-\frac{b^{2}}{a^{2}}$ and $(k^{\prime})^{2}=1-k^{2}$. $\mathbb{K}(k)$ and $\mathbb{E}(k)$ are complete elliptic integrals of the first and second kinds, respectively. The quantum corrections to the classical periods can be evaluated via the Picard-Fuchs technique as shown in (8). For the potential in (36), the coefficients $B_{k}^{(n)}$ depend only on the parameters $u_{1}$, $E$, and $m$. They can be determined systematically, with the first few orders given by
@@
These coefficients remain unchanged under wall-crossing. We further find that all coefficients $B_{k}^{(2n+1)}$ vanish for $n>0$, i.e., $\Pi_{\gamma}^{(2n+1)}=0$. This is a distinctive property of the cubic superpotential, which is closely related to the simplification of the corresponding $\mathbb{Z}_{4}$-extended TBA equations, as discussed below.
The marginal stability wall on the moduli parameters space $(u_{1},E)$ is located at
@@
In Fig. 1, the wall on the $E$ and $u_{1}$ moduli space is shown, respectively.
@@
To study the wall-crossing, we will break the parity symmetry of the potential $Q_{0}(x)$, which can be done by adding odd terms in $x$ to $Q_{0}(x)$. We will follow the path presented in Fig. 2 from the minimal chamber to the maximal chamber. The first chamber is the region of the moduli space of the potential corresponding to $|\phi_{2}-\phi_{1}|<\frac{\pi}{2}$ and $|\phi_{2}-\phi_{3}|<\frac{\pi}{2}$. The second chamber is determined by $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $|\phi_{2}-\phi_{3}|<\frac{\pi}{2}$ and $\phi_{12}-\phi_{3}<0$. The third one is $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $\frac{\pi}{2}<\phi_{2}-\phi_{3}<\frac{3\pi}{2}$ and $\phi_{12}-\phi_{3}<0$. The maximal chamber is determined by $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $\frac{\pi}{2}<\phi_{2}-\phi_{3}<\frac{3\pi}{2}$ and $\phi_{12}-\phi_{3}>0$. In each chamber, a new one-cycle will be added, which corresponds to a new Y-function in the TBA system. A similar process can be found in [6, 30].
@@
Figure 2: A path from the minimal to the maximal chamber by deforming the turning points of $Q_{0}(x)$ (red dots). The wavy lines denote the branch cut on the WKB curve. In each chamber, the one-cycles correspond to the Y-functions. In each process of wall-crossing, a new one-cycle (red cycle) and a new Y-function appear.
@@ 4.1 TBA in the minimal chamber
We thus obtain two independent $D_{3}$-type TBA-systems in [18, 30]. These $D_{3}$-type TBA systems also appear as the GMN TBA systems of $(A_{1},D_{3})$-type AD theory in the minimal chamber [20]. As observed in [30], the TBA systems (48) are only valid in the region $|m|\leq 1/2$. Beyond this region, one has to pick up the contribution of the pole at $1+\widehat{Y}=0$. For simplicity, we will focus only on the region $|m|\leq 1/2$.
@@
which have been tested numerically similarly to Table 1.
@@ 4.2 TBA in the maximal chamber
After the third wall-crossing in Fig.2, we arrive at the maximal chamber, where $4\times 6$ TBA equations are introduced; see (95) in Appendix A. Compared with those in the minimal chamber, three cycles $\gamma_{12}$, $\gamma_{32}$ and $\gamma_{312}$ have been added, which correspond to three types of new Y-functions $Y_{a,12},Y_{a,32}$ and $Y_{a,312}$, respectively, whose masses are given by
@@
The TBA equations (95) thus reduce to two independent TBA-systems (97). As in the case of the minimal chamber, we introduce
@@ 4.3 TBA at the monomial point
which also imposes constraints on the classical and next order correction of WKB periods. In conclusion, the monomial potential point in the maximal chamber is described by the two sets of $D_{3}$-type TBA equations, which are regarded as a $\mathbb{Z}_{4}$-extension of $A_{3}/\mathbb{Z}_{2}$ TBA equations [4].
@@ 5 Conclusion and discussion
In this paper, we have studied the exact WKB periods in deformed supersymmetric quantum mechanics. The exact WKB periods for any potential are described by the ${\mathbb{Z}}_{4}$-extended TBA equations, for which we have presented the general procedure of wall-crossing. In the case of cubic superpotential, the Schrödinger equation features a double-well potential with an additional $\hbar$-correction. When all turning points are real and distinct, the $\mathbb{Z}_{4}$ TBA equations decouple into two independent $D_{3}$-type TBA equations[30]. These coincide with the GMN TBA equations of $(A_{1},D_{3})$-type AD theory in the minimal chamber. After wall-crossing, the TBA equations reduce to two independent TBA systems as well, each matching the form of the GMN TBA equations of $(A_{1},D_{3})$-type AD theory in the maximal chamber. At the monomial point, where $W^{\prime}(x)^{2}=x^{4}$, the Schrödinger equation takes the same form as the monomial quartic potential but with an additional $\hbar$-correction in the potential. The TBA equations for the monomial point also separate into two independent $D_{3}$-type TBA systems [42], which provide a $\mathbb{Z}_{4}$-extension of the $A_{3}/\mathbb{Z}_{2}$ TBA equations.
@@
In the present work, we have considered the deformation parameter $m$ in the region $|m|\leq{1\over 2}$. Around $m=\pm 1/2$, the poles at $1+Y=0$ start to contribute, which makes the TBA equations unstable. It is important to consider the analytic continuation of TBA equations to cross $m=\pm 1/2$, which is related to the excited state in the integrable model. See [30] for the case of the minimal chamber. In our paper, we have mainly focused on the cubic superpotential. It is also interesting to consider the TBA in the maximal chamber for a more general superpotential. Imposing the symmetry at the monomial point, it is supposed to obtain a deformed TBA system of $A_{2N-1}/\mathbb{Z}_{2}$-type. Our results indicate a profound connection between the deformed supersymmetric quantum mechanics with cubic superpotential and the $(A_{1},D_{3})$-type AD theories, whose interpretation from the SW theory is an interesting open question. Moreover, the physical interpretation of the $\mathbb{Z}_{4}$-extension in the TBA equations within the framework of the integrable model would be interesting. The integrable model is characterized by the TBA equations. Two distinct $D_{3}$ TBA equations–given in (48) and (82)– are connected through the wall-crossing process. It would be valuable to explore how the corresponding integrable models are related, which is supposed to be a new deformation in the integrable model. It is also worth noting that the effective central charge is invariant in the process of wall-crossing. In the case of massless TBA in the minimal chamber [30], the effective central charge is found to relate to the supersymmetric index in the two-dimensional ${\cal N}=2$ theories [46]. It would be interesting to see the interpretation of the effective central charge from the ${\cal N}=2$ theories or the framework of the wall-crossing in the gauge theory.
@@ Appendix A $\mathbb{Z}_{4}$-extended TBA equations in the maximal chamber
For the cubic superpotential, we find the relations (67), which leads to the identifications (68). The TBA equations (95) thus reduce to two independent TBA-systems

### v2

@@ 4 TBA for the cubic superpotential
In this section, we illustrate our general procedure of wall crossing with the case of cubic superpotential, where the Schrödinger equation (1) becomes
@@
where $k^{2}=1-\frac{b^{2}}{a^{2}}$ and $(k^{\prime})^{2}=1-k^{2}$. $\mathbb{K}(k)$ and $\mathbb{E}(k)$ are complete elliptic integrals of the first and second kinds, respectively. The quantum corrections to the classical periods can be evaluated via the Picard–Fuchs technique as shown in (8). For the potential in (35), the coefficients $B_{k}^{(n)}$ depend only on the parameters $u_{1}$, $E$, and $m$. They can be determined systematically, with the first few orders given by
@@
These coefficients remain unchanged under wall crossing. We further find that all coefficients $B_{k}^{(2n+1)}$ vanish for $n>0$, i.e., $\Pi_{\gamma}^{(2n+1)}=0$. This is a distinctive property of the cubic superpotential, which is closely related to the simplification of the corresponding $\mathbb{Z}_{4}$-extended TBA equations, as discussed below.
The marginal stability wall on the moduli parameter space $(u_{1},E)$ is located at
@@
In Fig. 1, the walls on the $E$ and $u_{1}$ moduli spaces are shown, respectively.
@@
To study the wall crossing, we will break the parity symmetry of the potential $Q_{0}(x)$, which can be done by adding odd terms in $x$ to $Q_{0}(x)$. We will follow the path presented in Fig. 2 from the minimal chamber to the maximal chamber. The first chamber is the region of the moduli space of the potential corresponding to $|\phi_{2}-\phi_{1}|<\frac{\pi}{2}$ and $|\phi_{2}-\phi_{3}|<\frac{\pi}{2}$. The second chamber is determined by $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $|\phi_{2}-\phi_{3}|<\frac{\pi}{2}$ and $\phi_{12}-\phi_{3}<0$. The third one is $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $\frac{\pi}{2}<\phi_{2}-\phi_{3}<\frac{3\pi}{2}$ and $\phi_{12}-\phi_{3}<0$. The maximal chamber is determined by $\frac{\pi}{2}<\phi_{2}-\phi_{1}<\frac{3\pi}{2}$, $\frac{\pi}{2}<\phi_{2}-\phi_{3}<\frac{3\pi}{2}$ and $\phi_{12}-\phi_{3}>0$. In each chamber, a new one-cycle will be added, which corresponds to a new Y-function in the TBA system. A similar process can be found in [6, 30].
@@
Figure 2: A path from the minimal to the maximal chamber by deforming the turning points of $Q_{0}(x)$ (red dots). The wavy lines denote the branch cut on the WKB curve. In each chamber, the one-cycles correspond to the Y-functions. In each process of wall crossing, a new one-cycle (red cycle) and a new Y-function appear.
@@ 4.1 TBA in the minimal chamber
We thus obtain two independent $D_{3}$-type TBA systems in [18, 30]. These $D_{3}$-type TBA systems also appear as the GMN TBA systems of $(A_{1},D_{3})$-type AD theory in the minimal chamber [20]. As observed in [30], the TBA systems (47) are only valid in the region $|m|\leq 1/2$. Beyond this region, one has to pick up the contribution of the pole at $1+\widehat{Y}=0$. For simplicity, we will focus only on the region $|m|\leq 1/2$.
@@
which have been tested numerically, similarly to Table 1.
@@ 4.2 TBA in the maximal chamber
After the third wall crossing in Fig.2, we arrive at the maximal chamber, where $4\times 6$ TBA equations are introduced—see (94) in Appendix A. Compared with those in the minimal chamber, three cycles $\gamma_{12}$, $\gamma_{32}$ and $\gamma_{312}$ have been added, which correspond to three types of new Y-functions $Y_{a,12},Y_{a,32}$ and $Y_{a,312}$, respectively, whose masses are given by
@@
The TBA equations (94) thus reduce to two independent TBA systems (96). As in the case of the minimal chamber, we introduce
@@ 4.3 TBA at the monomial point
which also imposes constraints on the classical and next order correction of WKB periods. In conclusion, the monomial potential point in the maximal chamber is described by the two sets of $D_{3}$-type TBA equations, which are regarded as a $\mathbb{Z}_{4}$ extension of $A_{3}/\mathbb{Z}_{2}$ TBA equations [4].
@@ 5 Conclusion and discussion
In this paper, we have studied the exact WKB periods in deformed supersymmetric quantum mechanics. The exact WKB periods for any potential are described by the ${\mathbb{Z}}_{4}$-extended TBA equations, for which we have presented the general procedure of wall crossing. In the case of cubic superpotential, the Schrödinger equation features a double-well potential with an additional $\hbar$ correction. When all turning points are real and distinct, the $\mathbb{Z}_{4}$ TBA equations decouple into two independent $D_{3}$-type TBA equations[30]. These coincide with the GMN TBA equations of $(A_{1},D_{3})$-type AD theory in the minimal chamber. After wall crossing, the TBA equations reduce to two independent TBA systems as well, each matching the form of the GMN TBA equations of $(A_{1},D_{3})$-type AD theory in the maximal chamber. At the monomial point, where $W^{\prime}(x)^{2}=x^{4}$, the Schrödinger equation takes the same form as the monomial quartic potential but with an additional $\hbar$ correction in the potential. The TBA equations for the monomial point also separate into two independent $D_{3}$-type TBA systems [42], which provide a $\mathbb{Z}_{4}$ extension of the $A_{3}/\mathbb{Z}_{2}$ TBA equations.
@@
In the present work, we have considered the deformation parameter $m$ in the region $|m|\leq{1\over 2}$. Around $m=\pm 1/2$, the poles at $1+Y=0$ start to contribute, which makes the TBA equations unstable. It is important to consider the analytic continuation of TBA equations to cross $m=\pm 1/2$, which is related to the excited state in the integrable model. See [30] for the case of the minimal chamber. In our paper, we have mainly focused on the cubic superpotential. It is also interesting to consider the TBA in the maximal chamber for a more general superpotential. Imposing the symmetry at the monomial point, it is supposed to obtain a deformed TBA system of $A_{2N-1}/\mathbb{Z}_{2}$-type. Our results indicate a profound connection between the deformed supersymmetric quantum mechanics with cubic superpotential and the $(A_{1},D_{3})$-type AD theories, whose interpretation from the SW theory is an interesting open question. Moreover, the physical interpretation of the $\mathbb{Z}_{4}$ extension in the TBA equations within the framework of the integrable model would be interesting. The integrable model is characterized by the TBA equations. Two distinct $D_{3}$ TBA equations—given in (47) and (81)— are connected through the wall crossing process. It would be valuable to explore how the corresponding integrable models are related, which is supposed to be a new deformation in the integrable model. It is also worth noting that the effective central charge is invariant in the process of wall crossing. In the case of massless TBA in the minimal chamber [30], the effective central charge is found to relate to the supersymmetric index in the two-dimensional ${\cal N}=2$ theories [46]. It would be interesting to see the interpretation of the effective central charge from the ${\cal N}=2$ theories or the framework of the wall crossing in the gauge theory.
@@ Appendix A $\mathbb{Z}_{4}$-extended TBA equations in the maximal chamber
For the cubic superpotential, we find the relations (66), which lead to the identifications (67). The TBA equations (94) thus reduce to two independent TBA systems